OpenStudy (abbie):

What is the geometric mean of the pair of numbers? 96 and 6

OpenStudy (anonymous):

do you have choices?

OpenStudy (anonymous):

id multiply the numbers and take the square root

OpenStudy (anonymous):

Let the goemetric mean of the givenb numbers be x Hence \[\huge X^2 = 96 \times 6=576\] \[\huge \rightarrow X = \sqrt{576} \rightarrow X = 24\] Hence Geometric mean of the given numbers is 24.